Ms. Rosse is a Head Nurse at TCH. Her patient wanted an update of her progress. The temperatures are normally distributed with mean 101.1F and the standard deviation is 1.8F. She reads a report that says that in 16 readings the patient’s mean temp was 102.2F. Find the probability that the mean patient’s temperature was at most 102.2F.
Answer by stanbon(75887)

Ms. Rosse is a Head Nurse at TCH. Her patient wanted an update of her progress. The temperatures are normally distributed with mean 101.1F and the standard deviation is 1.8F. She reads a report that says that in 16 readings the patient’s mean temp was 102.2F. Find the probability that the mean patient’s temperature was at most 102.2F.
-----
t(102.2) = (102.2-101.1)/[1.8/sqrt(16)] = 2.4444
-----
P(temp > 102.2) = P(t > 2.4444 when df=15) = 0.0137
===============
Cheers
Stan H.
===========